OpenStudy (anonymous):

how do you solve an equation like this –5 = s – 3 ?

OpenStudy (saifoo.khan):

take the variables on one side, while the constants on other.

OpenStudy (anonymous):

Move the 3 to the other side. -5 + 3= s-3 + 3 -2=s

OpenStudy (saifoo.khan):

s = -5 + 3 Add like terms. s = -2

OpenStudy (anonymous):

thank you!

OpenStudy (saifoo.khan):

No Problem!